# Principles for Optimizing Quantum Transduction in Piezo-Optomechanical Systems

### v2

@@ V Discussion: Optimizing the transduction efficiency:
Changing the bus waveguide coupling $\kappa_{ex,2}$ can be accomplished independent of the overall optical coupling of the cladding MRR, $\kappa_{1}$, and the intrinsic optical coupling $\kappa_{0,2}$ (describing optical losses) in the external MRR. Similarly, as shown by the dashed red curve in Fig. 6, the maximum transduction efficiency increases with increased waveguide coupling, albeit at a higher power.
@@
Taken together, we can see that by simultaneously increasing the optomechanical, electromechanical, and bus waveguide coupling rates, the maximum transduction efficiency (dot-dashed yellow curve in Fig. 6) more than doubles its nominal maximum value (solid blue curve in Fig. 6) for reasonable experimental parameters. For the system considered here, we point out that there is no one element of the transducer that needs to be made of a material with both a high optomechanical coupling, and a high electromechanical coupling, as these processes occur in different parts of the transducer, and may be optimized independently. In addition, if we decrease the intrinsic losses $\gamma_{0}$, $\kappa_{1}$, and $\kappa_{0,2}$, the transduction efficiency can be increased to above $97\%$ (solid black curve in Fig. 6).
@@
It is important to note that contributions to the noise arising from sources other than the microwave signal have been discussed in [16], and are beyond the scope of what is discussed here. However, we point out that where the interactions in the transduction process are considered unitary, the sum of the squares of the overall coupling coefficients from each possible input mode to the output mode $\tilde{a}_{out}$ must add to unity, so that at higher transduction efficiency, we see a suppression of the conversion of noise photons into those at the transduced frequency. The amount of photons at the transduced frequency due to sources other than the intended input microwave photon will depend on the conditions of the experiments (e.g., thermal noise photons depending on temperature, power instabilities leading to relative intensity noise in the pump, etc). Careful consideration of noise becomes an issue when evaluating the transduction efficiency at the single-photon level, where, say, a cryogenic environment would be required to minimize the number of thermal microwave photons existing in the system that would in principle also be transduced. As current SC qubits operate at cryogenic temperatures, this is not a difficult condition to satisfy in realistic quantum transducers.
@@
V.1 Transduction efficiency vs bandwidth
Where our analysis has shown that high transduction efficiencies are achievable at a given microwave frequency, when that frequency is degenerate with the mechanical resonance frequency, it is equally important to consider over what range of frequencies these high efficiencies may be maintained. In Fig. 7, the transduction efficiency is plotted as a function of the input microwave frequency detuning relative to the mechanical resonance, $\omega-\omega_{m}$. At each frequency, the driving laser power is chosen as to maximize the transduction frequency on resonance.
@@ V.1 Transduction efficiency vs bandwidth
A surprising feature is that the peaks of these transduction spectra are not all on resonance, but have a shift that varies with the coupling constants (See Table 3 for shifts in each case). Ordinarily, this would be indicative of an error, but note that in our earlier analysis, we had neglected the frequency shift of the mechanical resonance due to the optomechanical interaction when linearizing the optomechanical hamiltonian. In the nominal case, this shift is small compared to the overall transduction bandwidth, but it grows as the coupling rates are increased. A related phenomenon is the resonance frequency shift of the optical cavity, which also increases as the coupling rates are increased, as seen in (25), but more analysis is needed to fully characterize how the shift varies with the different coupling parameters.
@@
nominal 0.0171069 1.73992
$5\kappa_{ex,2}$ 0.079846 1.54554
$5g_{EM}$ (broad peak) 12.3435
$5g_{EM},5\kappa_{ex,2},10G$ 1.22048 22.2126
$5g_{EM},5\kappa_{ex,2},10G,0.1\gamma_{0},0.1\kappa_{1}$ 1.20668 21.3662
@@
V.2 Proposed solution to frequency separation challenges by optomechanical four-wave mixing
As currently conceived, the piezo-optomechanical transducer may achieve high efficiency, but suffers from the substantial technical challenges of separating out the single photons of transduced light at telecom frequencies from a classically bright laser field in the same spatial mode only a few GHz away in frequency  [^7] Note: At $1550$nm, ±$\pm 1$GHz is ±$\pm 0.008$nm.. This contains three filtering challenges in one: first, to filter the pump light entering the transducer (or else have a narrow enough bandwidth pump) so that almost no pump photons exist at the transduction frequency; second, to separate the transduced photons generated from the remaining pump light; and third, to somehow mitigate the generation of photons at the transduced frequency due to spontaneous four-wave mixing with a high-power pump producing photon pairs (albeit at a suppressed rate by the two-ring design) that are still well-phase matched due to the small frequency separation. Two out of three of these filtering challenges do not apply in the case of measuring the transduction efficiency at classical microwave intensities, which have been accomplished using a self-calibrated heterodyne detection method [17].
Another filtering scheme involves sending the output optical signal from the transducer through cascaded high-finesse fiber Fabry-Perot filters, which filter out the pump photons before sending the transduced photons to a single photon detector [39]. While this scheme successfully filters out pump photons, it is complex and expensive, and requires routing optical signals out of the cryrogenic environment for filtering. However, this need not be a liability when quantum networking would require routing optical signals out of the cryogenic environment anyway.
Figure 8: Diagram of concatenated frequency filtering using integrated optics. In the absence of retroreflections, each time the light is transmitted through an MRR to the next waveguide, its spectrum is effectively multiplied by a Lorentzian function.
There are many means of frequency separation that work well for large differences in frequencies, but the interferometric methods used to separate small differences in frequencies do not yet have a high enough extinction ratio to be used in this system, particularly in integrated optics. Some of this is fundamental, as a lorentzian-shaped pump spectrum would need to have an extremely narrow linewidth because its tails decay only as the second power of the frequency difference, instead of exponentially fast. At $1550$nm, a $100$mW laser has on average $7.80\times 10^{17}$ photons per second, which means we would need an extinction ratio better than $179dB$ for this light to be attenuated to less than one photon per second, so that the transduced light begins to become significant by comparison. If we consider a Lorentzian filter spectrum centered at 1550nm that needs to decay by more than 179dB with a detuning of only 3 GHz, the linewidth of said pump would need to be incredibly narrow at about $6.7$Hz. However, this does not account for the possibility of cascading these filters. Using a double-bus micro-ring resonator, the fraction of the input light that travels from the first bus, around the MRR, and exits through the second bus via the “drop” port (on the same side as the input) constitutes an approximately Lorentzian transmission spectrum. If we assume that any back-reflection due to scattering is negligible, then concatenating these filters amounts to adding a new ring some distance along the second bus waveguide that will couple also to a third bus, and then another ring some distance along the third bus waveguide that will couple to a fourth bus waveguide (and so on, so that the light zigs and zags its way through the rings and waveguides (see Fig. 8 for diagram)). Without the complication of back-reflections, we can assume that concatenating $N$ identical filters in this way raises the transmission spectrum to the $N^{\text{th}}$ power, greatly broadening the maximum transmission linewidth of each filter without compromising the required attenuation $3$ GHz away. In fact, the corresponding linewidth requirement for two cascaded filters would be about $200$kHz per filter; for three filters would be about $6.23$ MHz per filter, and for five cascaded filters would be about $97.3$ MHz per filter, which are much more realizable, assuming filters cascaded this way can be kept stable. Note: When back-scattering is small, but cannot be neglected, we may still use this approach so long as the effective round trip loops made possible with multiple backreflections are not near resonance with the frequency of interest. See Appendix E for details.
For an alternative design of an optical filter, we can consider recent developments in high extinction CROW (Coupled Resonator Optical Waveguide) filters [40] in comparison with the cascaded filter design we have proposed. In Ref. [41], they use a pair of concatenated second-order CROW filters to achieve an extinction ratio in excess of $90$dB without incurring substantial losses at the wavelength of interest. Their approach of using two second-order CROW filters instead of one fourth-order CROW filter was for its relative ease of fabrication with respect to manufacturing tolerances. Where they were using these filters for pump removal from photons generated by SFWM in silicon waveguides separated from that pump by about $400$ GHz (a much wider separation than the $3$ GHz separation we seek to achieve), this class of filters still holds promise as manufacturing capabilities improve.
Instead of addressing this formidable technical challenge of high-extinction filtering over a narrow transition band directly (which would still not address the spurious photons generated from the pump via spontaneous four-wave mixing), one can instead reconsider and engineer the transducer’s optomechanical interaction to the same desired end.
The form of the optomechanical interaction studied thus far is a three-wave mixing interaction that is second-order in the electromagnetic field, and first-order in the acoustic field. However, if one could achieve a four-wave mixing interaction that is third-order in the electromagnetic field, and still first-order in the acoustic field, one could employ two classically bright pumps at approximately half the frequency (twice the wavelength) of the transduced light, allowing them to both be widely separated (and easy to filter) from the transduced light, as well as being of a low enough energy that neither Raman-Stokes scattering nor spurious all-optical SFWM from the pump would be a significant source of noise. Moreover, the conditions under which this four-wave-mixing optomechanical interaction would be phase-matched would actually preclude the generation of light at twice the pump frequency via $\chi^{(2)}$ sum frequency generation, since both processes cannot be simultaneously phase matched without exceptional consideration. In short, there should be no confounding processes producing light also at the transduced frequency when using optomechanical four-wave mixing.
Note that in principle, a purely optical four-wave mixing ($\chi^{(3)}$) interaction could be used for frequency conversion with the advantage of easy filtering of pump from transduced light, but it remains unfeasible as multiple-competing $\chi^{(3)}$ processes become efficient under the same conditions that would optimize this frequency conversion process. The four-wave mixing optomechanical interaction does not suffer from this same disadvantage.
To achieve this four-wave mixing optomechanical interaction, we consider higher-order versions of photoelasticity and electrostriction, which are thermodynamically connected in the ordinary optomechancial interaction. Let us define the cubic electrostriction $q_{ijk\ell m}$
$q_{ijk\ell m}\equiv{\color[rgb]{0,0,0}\epsilon_{0}^{2}}\frac{\partial^{3}X_{\ell m}}{\partial D_{i}\partial D_{j}\partial D_{k}}.$ (69)
As discussed in Appendix C, one can use the equality of mixed partial derivatives of the Helmholtz free energy to show that this is proportional to the response of the second-order nonlinear inverse permittivity $\eta_{ijk}^{(2)}$, to strain (i.e., a second-order photoelasticity):
$q_{ijk\ell m}=2{\color[rgb]{0,0,0}\epsilon_{0}^{2}}\frac{\partial\eta^{(2)}_{ijk}}{\partial x_{\ell m}}.$ (70)
As defined, the second-order photoelasticity modifies $\eta_{ijk}^{(2)}$ similarly to how ordinary photoelasticity modifies $\eta_{ij}^{(1)}$:
$\eta_{ijk}^{(2)}(x_{\ell m})\approx\eta_{ijk}^{(2)}(0)+\frac{1}{2{\color[rgb]{0,0,0}\epsilon_{0}^{2}}}q_{ijk\ell m}x_{\ell m}$ (71)
In this way, we understand that sound can be used to modulate $\eta_{ijk}^{(2)}$ in a manner similar to periodic poling, though the accompanying modulation of the first-order susceptibility at a smaller magnitude makes this function simultaneously as an acousto-optic modulator or a diffraction grating depending on the poling period desired.
In the same way that we obtained the ordinary optomechanical interaction hamiltonian (57) by considering corrections to the electric displacement field, we can obtain the following four-wave mixing optomechanical interaction hamiltonian:
$\hat{H}_{int}=\frac{1}{6\epsilon_{0}^{2}}\int d^{3}r\left(q_{ijk\ell m}^{(eff)}\hat{D}_{i}\hat{D}_{j}\hat{D}_{k}\hat{x}_{\ell m}\right)$ (72)
If we consider the two pump fields to be the same pump field, we can simplify this interaction further, linearizing it to create the optomechanical interaction hamiltonian of the form:
$\hat{H}_{int}\approx\hbar g_{eff}\left[\left(\bar{a}^{*2}\!+\!2|\bar{a}|^{2}\right)\delta\hat{a}+\left(\bar{a}^{2}\!+\!2|\bar{a}|^{2}\right)\delta\hat{a}^{\dagger}\right]\!\!\left(\!\hat{b}+\hat{b}^{\dagger}\!\right)$ (73)
where here, terms second-order and higher in $\delta\hat{a}$ have been approximated as zero, as in the standard linearization procedure. Ultimately, we can see that this four-wave-mixing optomechanical interaction Hamiltonian is very similar in form to the ordinary optomechanical interaction (61), except that the coupling constant depends on stress-induced nonlinear susceptibility, and that the overall coupling constant is quadratic in the field instead of linear in the field. Where the stress-induced nonlinearity may be small, the overall interaction may still be efficient because it grows more rapidly with the applied pump field than it would in the ordinary optomechanical interaction.
To make this four-wave mixing interaction efficient, we would need materials that exhibit a high degree of second-order photoelasticity, as well as materials with the right acoustic and optical dispersion to maximize the overlap integral (i..e, optomechanical phase matching). The second-order photoelasticity does not yet appear to be tabulated in the literature, but we can with some approximation determine a formula for it using known material parameters.
Consider Miller’s rule in nonlinear optics [42, 30]. In particular, it provides a means to estimate the change in the second-order optical susceptibility for a particular set of frequencies, as it is proportional to the product of the first-order optical susceptibilities at the respective frequencies being considered. In terms of inverse optical susceptibilities, Miller’s rule takes the form:
$\eta^{(2)}_{ijk}(\omega_{3}=\omega_{1}+\omega_{2})=-Q\epsilon_{0}\prod_{n=1,2,3}\left(\frac{1}{\epsilon_{0}}-\eta^{(1)}_{nn}(\omega_{n})\right)$ (74)
where $Q$ is a constant of proportionality. Because Miller’s rule can be used to transpose between different values of the nonlinear susceptibility given perturbations to the refractive index (which changes the first-order susceptibility), we can expect that materials that simultaneously have a large second-order nonlinearity, and a large photoelasticity (e.g., barium titanate) will have a correspondingly large degree of second-order photoelasticity. Indeed, if we assume the constant of proportionality $Q$ is unchanging for each material (and that we are in a coordinate frame that diagonalizes $\eta_{ij}^{(1)}$), then we may obtain the approximate empirical formula for the second-order photoelasticity without needing to know the value of $Q$:
$q_{ijk\ell m}=2\epsilon_{0}^{2}\frac{\partial\eta_{k\ell m}^{(2)}}{\partial x_{ij}}\approx-2\epsilon_{0}^{2}\eta^{(2)}_{k\ell m}\sum_{n=1,2,3}\frac{p_{nnij}(\omega_{n})}{\frac{1}{\epsilon_{0}}-\eta_{nn}^{(1)}(\omega_{n})}$ (75)
This is accomplished by differentiating our expression for $\eta_{ijk}^{(2)}$ (74) with respect to strain, and eliminating $Q$ in the resulting expression by substituting its value as determined by (74).
Our expression for the second-order photoelasticity may be simplified and approximated to:
$q^{(eff)}\approx-\frac{4d_{eff}}{n^{2}(\omega_{1})n^{2}(\omega_{2})n^{2}(\omega_{3})}\sum_{n=1,2,3}\frac{p_{nnij}(\omega_{n})}{1-\frac{1}{n^{2}(\omega_{n})}}$ (76)
where the indices of refraction at the given frequencies are taken along the appropriate polarizations, and $d_{eff}$ is half the effective value of the second-order nonlinear optical susceptibility more commonly tabulated in the literature (and accounting for quasi-phase matching factors as necessary). For the transduction process under consideration, we let $(\omega_{1}=\omega_{p1},\omega_{2}=\omega_{p2},\omega_{3}=\omega_{t}=\omega_{p1}+\omega_{p2}+\omega_{m})$.
With this formula, we can enter constants to see what a typical value for the second-order photoelasticity might be. As an example, for Barium titanate, where ($n(1550nm)\approx 2.3$, $n(775nm)\approx 2.4$, $d_{eff}\approx 10pm/V$, $p_{1133}\approx 0.2$, $p_{2233}\approx 0.2$,$p_{3333}\approx 0.77$), this gives an effective second-order photoelasticity $q^{(eff)}\approx 0.35pm/V$. Combining this with the constants for the operators in the interaction gives us an approximate formula for this new four-wave optomechanical coupling (73):
$g_{eff}$ $\approx\frac{\hbar q_{eff}}{24\epsilon_{0}^{2}}\sqrt{\frac{\omega_{p1}\omega_{p2}\omega_{t}}{\omega_{m}}}$
×$\times\sqrt{\frac{1}{\eta_{eff}(\omega_{p1})\eta_{eff}(\omega_{p2})\eta_{eff}(\omega_{t})\rho_{eff}(\omega_{m})}}$
×$\times\sqrt{\frac{1}{V_{eff}^{(m)}V_{eff}^{(p1)}V_{eff}^{(p2)}V_{eff}^{(t)}}}$
×$\times\int\!\!d^{3}r\,\mathcal{E}_{p1}(\vec{r})\mathcal{E}_{p2}(\vec{r})\mathcal{E}^{*}_{t}(\vec{r})\frac{\partial w_{i}(\vec{r})}{\partial r_{j}}$ (77)
With this we have the full dependence of optomechanical four-wave mixing. While we know that $q_{eff}$ is large for materials for a large second-order nonlinearity and ordinary photoelasticity, we can also understand that a material with refractive index as close to unity as possible (all other factors constant) may have a greatly enhanced efficiency for this matter of optomechanical coupling. Moreover, we see the overlap integral for four fields appears, which provides the same manner of optomechanical phase matching (albeit for four wave mixing).
Assuming an approximately rectangular geometry (similar to a ring resonator under conformal transformation as long as it’s within resonance), we can give this optomechanical phase matching by the expression:
[…]
where:
$\Delta k_{z}$ $=k_{tz}-k_{p1z}-k_{p2z}-k_{mz}-k_{\Lambda z}$
$=\frac{n(\omega_{t})\omega_{t}}{c}-\frac{n(\omega_{p1})\omega_{p1}}{c}-\frac{n(\omega_{p2})\omega_{t2}}{c}-\frac{2\pi}{\Lambda}$ (79)
Note here the absence of the acoustic momentum in the $z$ direction. In the geometries we consider, the acoustic momentum is perpendicular to the direction of the light propagation. Assuming the acoustic wavelength is much longer than the (transverse) dimensions of the waveguide, we may neglect the corresponding transverse phase matching integral, as the transverse phase matching bandwidth will be so broadband that other factors in the overlap integral will dominate the transverse behavior. In this way, we see that the poling momentum $k_{\lambda}$ would be what ever is necessary to compensate for the momentum offset between the transduced photon, and the two pump photons. This at first appears tricky, as one has to match the poling period to the circumference of the micro-ring resonator. However, with a ring radius of only 10 microns, and typical poling periods being tens of microns, one will have to adjust the ring radius by a small amount to ensure an integer number of poling periods in one round trip.
Besides selecting materials with a strong second-order photoelasticity, the system would also have to be designed to take into account how the disparate wavelengths of the pump and the transduced light will couple into/out of the waveguide, and how to keep them in the same transverse spatial mode to maximize the overlap. This is nontrivial because the wavelengths of the pump photons and transduced light will be far enough apart that pump light may not be single-mode in the same waveguide where the transduced light would be.
Note that this interaction may work well for a wider variety of materials than those with a large second-order nonlinear susceptibility because we only consider the stress-induced change to it rather than its nominal value. That said, centrosymmetric crystals, which have zero second-order susceptibility will likely also have a zero second-order photoelasticity under the small strains considered here  [^8] When viewed from the center of a centrosymmetric crystal’s unit cell, all strains at this scale are symmetric, and every atom’s perturbation from the strain is matched by an equal and opposite perturbation for the corresponding atom opposite the center of symmetry. Centrosymmetry remains unbroken under small strain for these systems, and under these conditions will not have a stress-induced second-order nonlinearity., (though second-order photoelasticity exists for particularly large strains [44] such as those that can be frozen-in due to anisotropies in the manufacturing process).
In addition, going from the three-wave mixing of conventional optomechanical coupling to the four wave mixing of this proposed optomechanical coupling adds the complexities of additional processes that can change how light propagates through the material (e.g., self-phase modulation, among others). In particular, equation (73) shows that in in this higher-order optomechanical interaction, the phase of the pump itself affects the magnitude of the optomechanical coupling (where in the conventional case it depends only on magnitude), and so must be managed in addition to having an overall high intensity to optimize transduction.
V.3 Relative transduction efficiency using optomechanical four-wave mixing
At this point it is worth considering that the strain from a single photon is generally negligibly small compared to any classical vibration. Indeed, if we consider a classical sound source with power of about $10^{-12}W$ (or about $0dB$), for phonons of $3GHz$, this would correspond to having a rate of of $4.5\times 10^{12}$ phonons per second. With zero dB as approximately the scale of perceptible hearing, and the power of a sound wave being proportional to the square of its strain amplitude, this means that the strain associated to a single phonon must be about six orders of magnitude below what is considered barely perceptible in hearing. However, with so small a strain, we may still affect near unity transduction efficiency with a sufficiently powerful pump laser. The question we must answer here is whether it is reasonable to expect comparable efficiencies when using optomechanical four-wave mixing.
To estimate this, we consider how the optomechanical coupling constant $g_{OM}$ when using optomechanical four-wave mixing may be used in the same way as the ordinary optomechanical three-wave mixing in the transducer studied in this paper [16, 17]. In nonlinear optics, a $\chi^{3}$ nonlinear medium with a static DC electric field acting on it may cause the material to act like a $\chi^{2}$ medium. The nonlinear-optical hamiltonian for the $\chi^{3}$ material may be given by:
$\hat{H}_{NL}\!\!$ $=\!\!\frac{1}{4}\!\!\int\!\!\!d^{3}r\big{(}\big{[}\eta_{ijk\ell}^{(3)}(\vec{r})\hat{D}_{i}^{-}\!(\vec{r}\!,\!t)\big{]}\hat{D}_{j}^{-}\!(\vec{r}\!,\!t)\hat{D}_{k}^{+}\!(\vec{r}\!,\!t)\hat{D}_{\ell}^{+}\!(\vec{r}\!,\!t)\big{)}$
$\qquad+H\!.c..$ (80)
When one of the fields is a static DC electric field, the product in brackets can be consolidated into an effective second-order (inverse) nonlinearity:
$\frac{3}{4}\eta_{ijk\ell}^{(3)}(\vec{r})\hat{D}_{i}^{-}\!(\vec{r}\!,\!t)=\eta_{jk\ell}^{(2),eff}(\vec{r})$ (81)
For a sufficiently strong electric field, one may induce an effective second-order nonlinearity in silicon (which ordinarily has none) to have a $d_{eff}$ comparable to conventional second-order crystals. From this, we can understand in optomechanical four-wave mixing, when one of the fields may be considered constant, we may make the effective approximation:
$\frac{q_{ijk\ell m}^{eff}}{3\epsilon_{0}}\hat{D}_{i}\approx\frac{q_{ijk\ell m}^{eff}}{3}\epsilon_{R}\hat{E}_{i}\approx p_{ijk\ell}^{(eff)}$ (82)
For our example with barium titanate, we had a $q^{(eff)}\approx 0.35\times 10^{-12}m/V$, and $\epsilon_{R}\approx 2000$. In these units, $p_{ijk\ell}^{(eff)}\approx 2.3\times 10^{-10}\hat{E}_{i}$. Where the average electric field of a $1mW$ Gaussian beam of light propagating through a barium titanate waveguide with mode field diameter (MFD) of $700nm$ has an average peak electric field amplitude given by the relation:
$|E|\approx\sqrt{\frac{4P}{\pi\epsilon_{0}c\,n(MFD)^{2}}}$ (83)
which within an order of magnitude would be approximately $2.2\times 10^{7}$ (two orders of magnitude below its approximate breakdown field of $2\times 10^{9}V/m$ [45], (for comparison, the breakdown field in air is only about $3\times 10^{6}$/m)), giving an effective photoelasticity $p^{(eff)}$of about $0.005$. While this is rather small compared to the nominal photoelasticity of $0.77$ for barium titanate, this value will continue to increase at higher pump powers, and the same design solutions that may optimize the ordinary optomechanical coupling will improve this manner of optomechancial coupling as well, and at a higher rate with some methods.
VI Conclusion

### v4

@@ IV.1.2 Optimal system parameters
Changing the bus waveguide coupling $\kappa_{ex,2}$ can be accomplished independent of the overall optical coupling of the cladding MRR, $\kappa_{1}$, and the intrinsic optical coupling $\kappa_{0,2}$ (describing optical losses) in the external MRR. Similarly, as shown by the dashed red curve in Fig. 4, the maximum transduction efficiency increases with increased waveguide coupling, albeit at a higher power. This can be understood because although increasing $\kappa_{ex,2}$ increases $\mathscr{F}_{2}$, while decreasing $\mathcal{C}_{12}$, the derivative of the transduction efficiency with respect to $\kappa_{ex,2}$ remains positive so long as $\mathscr{F}_{2}$ is below the threshold value $(1+\mathcal{C}_{12})/(2+\mathcal{C}_{12})$, which is above $99.99\%$ for the transducer studied here, making it a straightforward condition to satisfy.
Note: it may seem surprising that one needs higher pump power to recover the optimum transduction efficiency following increasing $\kappa_{ex,2}$ given that a decreasing $\mathcal{C}_{12}$ implies the optimum value of $\mathcal{C}_{O\!M}=\mathcal{C}_{12}+1$ decreases as well, so that $|\bar{a}_{1}|^{2}$ would actually need to be reduced to attain the maximum efficiency. This all is true, but increasing $\kappa_{ex,2}$ indirectly decreases $\mathcal{C}_{O\!M}$ because it decreases the cavity enhancement factor (5) between input pump power $|\bar{a}_{in}|^{2}$ and the intra-ring pump power $|\bar{a}_{1}|^{2}$.
By choosing materials with a strong photoelastic response, among other properties, one can also tune the optomechanical coupling rate, though this quantity only enters into the transduction efficiency formula via the driving laser amplitude  [^6] This is under the assumption that the optomechanical coupling is not a substantial source of loss via electrostriction contributing to $\kappa_{0,2}$.. Because of this, changing the optomechanical coupling does not change the maximum transduction efficiency, but increasing it does decrease the required laser power needed in order to achieve it.
It is interesting to consider that ultimately, the transduction efficiency asymptotically approaches zero at sufficiently high optical power. One interpretation of this effect would be when the coupling is too large, the transduced light may be generated either directly via transduction, or as a higher-order effect where part of the transduced light is back-converted and transduced again. These modes of transduced light may interfere destructively with one another in a manner similar to what goes on in optomechanically-induced transparency [46, 47].
@@
Taken together, we can see that by simultaneously increasing the optomechanical, electromechanical, and bus waveguide coupling rates (from their nominal values in [16]), the maximum transduction efficiency (dot-dashed yellow curve in Fig. 4) more than doubles its nominal maximum value (solid blue curve in Fig. 4) for reasonable experimental parameters. For the system considered here, we point out that there is no one element of the transducer that needs to be made of a material with both a high optomechanical coupling, and a high electromechanical coupling, as these processes occur in different parts of the transducer, and may be optimized independently. In addition, if we decrease the intrinsic losses $\gamma_{0}$, $\kappa_{1}$, and $\kappa_{0,2}$ we directly increase the maximum transduction efficiency by increasing $\mathscr{F}_{m}$, $\mathscr{F}_{2}$, and $\mathcal{C}_{12}$. As an example, the solid black curve in Fig. 4 shows the maximum transduction efficiency for the system studies here can be increased to above $97\%$ if the intrinsic losses can be reduced by an order of magnitude.
IV.1.3 Mitigation of noise at high efficiency
At high transduction efficiency, we must answer the question of whether or not the system will be overwhelmed by noise photons in the transduction band washing out the transduced quantum signal. Contributions to the noise arising from sources other than the microwave signal have been discussed comprehensively in [16], and will not be reproduced here. However, just as the transduced light generated from the input microwave light is computable as an amplitude, we may compute the amplitudes to produce light at the transduction frequency from every noise source considered in the transducer. To the extent that the interactions in the transduction process are considered unitary, the sum of the squares of the overall coupling coefficients from each possible input mode (input and noise) to the output mode $\tilde{a}_{\text{out}}$ must add to unity, so that at higher transduction efficiencies, we see a suppression of the conversion of noise photons into those at the transduced frequency.
@@ IV.1.3 Mitigation of noise at high efficiency
The amount of photons at the transduced frequency due to sources other than the intended input microwave photon will depend on the conditions of the experiments (e.g., thermal noise photons depending on temperature, power instabilities leading to relative intensity noise in the pump, etc). Careful consideration of noise becomes an issue when evaluating the transduction efficiency at the single-photon level, where, say, a cryogenic environment would be required to minimize the number of thermal microwave photons existing in the system that could in principle also be transduced. As current SC qubits operate at cryogenic temperatures, this is not a difficult condition to satisfy in realistic quantum transducers.
@@
IV.2 Transduction efficiency vs bandwidth
Where our analysis has shown that high transduction efficiencies are achievable at a given microwave frequency, when that frequency is degenerate with the mechanical resonance frequency, it is equally important to consider over what range of frequencies these high efficiencies may be maintained. In Fig. 6, the transduction efficiency is plotted as a function of the input microwave frequency detuning relative to the mechanical resonance, $\omega-\omega_{m}$. At each frequency, the driving laser power is chosen as to maximize the transduction efficiency on resonance.
@@ IV.2 Transduction efficiency vs bandwidth
A surprising feature is that the peaks of these transduction spectra are not all centered on same value, but appear to have a shift that varies with the coupling constants (See Table 3 for shifts in each case). Moreover, the plots in Fig. 6 are taken at the power level that maximizes the peak transduction efficiency on resonance (implicitly setting $\mathcal{C}_{O\!M}(\omega)=\mathcal{C}_{12}(\omega)+1$), which ought to imply that all the transduction spectra have their maximum at the same frequency. However, in our earlier analysis, we had neglected the frequency shift of the mechanical resonance due to the optomechanical interaction when linearizing the optomechanical hamiltonian. In the nominal case, this shift is small compared to the overall transduction bandwidth, but it grows as the coupling rates are increased. A related phenomenon is the resonance frequency shift of the optical cavity, which also increases as the coupling rates are increased, as seen in (6), but more analysis is needed to fully characterize how the shift varies with the different coupling parameters.
@@
nominal 0.0171 1.740
$5\kappa_{ex,2}$ 0.0799 1.546
$5g_{EM}$ (broad peak) 12.344
$5g_{EM},5\kappa_{ex,2},10G$ 1.220 22.213
$5g_{EM},5\kappa_{ex,2},10G,0.1\gamma_{0},0.1\kappa_{1}$ 1.207 21.366
@@
V Conclusion